OpenStudy (anonymous):

A flower bed is in the shape of a triangle with one side twice the length of the shortest side, and the third side is 29 feet more than the length of the shortest side. Find the dimensions if the perimeter is 193 feet.

OpenStudy (kkutie7):

We can start here. P=a+b+c a=shortest side c=a+29 P=193 b=2*a

OpenStudy (kkutie7):

from there we can assume: 193=a+b+c 193=a+(2*a)+c 193=a+(2*a)+(a+29)

OpenStudy (kkutie7):

193=a+2a+a+29 193=4a+29 164=4a 41=a

OpenStudy (kkutie7):

c=a+29 c=41+29 c=70

OpenStudy (kkutie7):

b=2*a b=2*41 b=82

OpenStudy (kkutie7):

there are your dimensions =)